# Breakdown of chiral anomaly and emergent phases in Weyl semimetals under orbital magnetic fields

### v1

@@ IV.1 Bloch-Hofstadter Hamiltonian
Translation symmetry can be restored when the flux is commensurate with the lattice. This is done by choosing an enlarged unit cell containing an integer number of flux quanta. In what follows we will choose commensurate flux $\phi_{B}/\phi_{0}=p/q$, where $q$ is a positive integer and without loss of generality we will set $p=1$ throughout this paper. The enlarged unit cell is then obtained by extending the original unit cell $q$ times along the $y$ direction. This gives a magnetic BZ with $\pi\leq k_{x}a\leq\pi$, $\pi/q\leq k_{y}a\leq\pi/q$, $\pi\leq k_{z}a\leq\pi$. Working with the magnetic BZ will allow us to compute topological invariants and track the evolution of the Fermi arc states in a straightforward manner. Although $k_{x}$ in the magnetic BZ lies in the range $(-\pi,\pi)$ the spectrum along $k_{x}$ is in fact periodic with a period of $2\pi/q$ [Bernevig2013]. Therefore, in the analysis below we restrict ourselves to $k_{x}\in(-\pi/q,\pi/q)$.
@@ IV.2 Phase diagram for $\gamma<0$
For the special case $\tilde{v}_{x}=v_{y}=1$, we can understand the transition between NI and LCI analytically by solving for the zeros of the Bloch-Hofstadter Hamiltonian in Eq. 9. These correspond to Weyl nodes in the magnetic BZ [Abdulla_Murthy_2022, Abdulla_2024], and their appearance is determined by the condition
@@
Figure 7: Energy gap as a function of separation between the Weyl nodes for sizable fluxes (a) $\phi_{B}/\phi_{0}=1/10$ and (b) $\phi_{B}/\phi_{0}=1/5$. See text for a detailed discussion of the sequence and structure of the transitions between the different gapped phases. (c) In a transition from an NI to an LCI the system goes through several closings and reopenings of the gap, as shown in Fig. 8. The red dots in (c) correspond to the values of $Q$ at which the energy bands in Fig. 8 are plotted.
@@
To understand the topological nature of the gapped phases, we calculate the $k_{x}$-dependent Chern number $C(k_{x})$ in the magnetic BZ. We find that for large $g$ (i.e. small $Q<Q^{*}$) $C(k_{x})=0$ for all $k_{x}$, corresponding to an NI. On the other hand, for small $g$ (i.e. large $Q>Q^{*}$), $C(k_{x})=1$ for all $k_{x}$, corresponding to an LCI phase. The transition between NI and LCI can be understood from the evolution of the solutions to Eq. (10). Let us fix the flux $1/q$ and look for solutions as $g$ (or equivalently $Q$) is varied. Starting from large $g$ (NI phase), when $g$ is decreased to $g_{2}$, the band gap closes and a Dirac point appears at the Gamma point of the magnetic BZ. Decreasing $g$ further leads to a splitting of the Dirac point into a pair of WNs that move away from each other towards the zone boundary. In this regime, solving in the slab geometry, a Fermi arc appears at the surface, and its length increases with the separation between the WNs. At a second critical value, $g=g_{1}$, the two WNs merge at the edge of the magnetic BZ $(\pm\pi/q,0,0)$ and disappear but the Fermi arc survives and spans the full magnetic BZ along $k_{x}$. Thus an LCI state emerges. This is demonstrated in Fig. 4.
@@ IV.3 Phase diagram for $\gamma>0$ and appearance of LCI
Figure 9: Evolution of zero energy surface states in the $k_{x}$-$k_{y}$ surface BZ in the transitions from an NI to an LCI state (see Fig. 7(c)). Figure (a) shows Fermi-arc states of the WSM state which appear in the transition NI to LCI. Chern numbers are calculated for a fixed $k_{x}$ value by integrating the Berry curvature over $k_{y}$ and $k_{z}$. (b) As $Q$ is increased, the Weyl nodes approach the edge of the magnetic BZ, merge, and get annihilated. At the end of this process, the corresponding Fermi arc spans the entire BZ along $k_{x}$. Thus an LCI state emerges. (c) A pair of Weyl nodes appear, accompanied by a new Fermi arc. This Fermi arc coexists with the surface states of the LCI bands. (d) As $Q$ is increased further, the Weyl nodes approach the edge of the magnetic BZ, merge at $(\pi/q,\pi/q)$, and get annihilated. Again, the corresponding Fermi arc now spans the entire BZ along $k_{x}$. Thus an LCI state emerges. In this state $C(k_{x})=0$ for all $k_{x}$.
@@ V Summary and Conclusion
An external orbital magnetic field, which is perpendicular to the direction of separation of Weyl cones of opposite chirality, introduces tunneling of electrons between the Weyl cones. This leads to a hybridization between the cones (note that such a hybridization is allowed since translation symmetry is broken by the presence of the orbital magnetic field) and an opening of a gap. Such a tunneling process requires overcoming the energy barrier separating the two Weyl nodes, leading to a nonperturbative dependence of the gap on the magnetic field. The induced gap is significant when the inverse magnetic length is comparable to or larger than the momentum-space separation between the Weyl nodes.
@@
We now turn to discuss briefly some of the experimental implications of our results. The magnetic field-induced gap opening in WSMs has direct consequences on the longitudinal magnetoresistance [Bednik_Syzranov_2020]. Such behavior has been already observed in experiments in the Weyl materials TaAs [Zhang_Jia_2017] and TaP [Ramshaw_McDonald_2018]. However, longitudinal magnetoresistance does not give a direct indication of the topology of the gapped phases. In contrast, Hall conductivity measurements provide a clearer distinction between the gapped phases, namely the normal insulator, LCI, and LCI phases. A time-reversal symmetry-breaking Weyl semimetal with two Weyl nodes separated by a momentum space distance $Q$ along $k_{x}$ exhibits a non-quantized anomalous Hall conductivity of $\sigma_{yz}=\frac{e^{2}}{2\pi h}Q$ per layer along the $x$-direction. Upon entering the LCI phase with applied magnetic field, this Hall conductivity becomes quantized at $\sigma_{yz}=\frac{e^{2}}{h}$ per layer. In contrast, transitions to either a normal insulator or the LCI phase result in a vanishing Hall conductivity. To further distinguish between normal insulator and LCI, one may rely on surface-sensitive probes, such as STM, that can detect the presence of the zero-energy modes arising from the closed Fermi arc states in LCI.
Finally, we comment on orbital magnetic fields that are not perfectly perpendicular to the separation of the Weyl nodes. The momentum parallel to the magnetic field, $k_{\parallel B}$, is a good quantum number. When the field has a component parallel to the separation between Weyl nodes, the value of $k_{\parallel B}$ differs at the two Weyl nodes, and they no longer hybridize directly. Hence, even a small misalignment of the field relative to the perpendicular direction can drastically suppress the gap opening [Bednik_Syzranov_2020]. However, this effect is mitigated in systems with nodal loops[Fang_Topological2015, Abdulla_SNS_2023, Abdulla_Internal_2025, Bian_Topological_2016, Bzdusek2016, Neupane_Observation_2016, Lou2018, Belopolski2019, Chen2021]. In such systems, magnetic fields generically lead to direct hybridization of nodal states, regardless of the magnetic field orientation. This will be discussed in detail elsewhere.

### v2

@@ IV.1 Bloch-Hofstadter Hamiltonian
Translation symmetry can be restored when the flux is commensurate with the lattice. This is done by choosing an enlarged unit cell containing an integer number of flux quanta. In what follows we will choose commensurate flux $\phi_{B}/\phi_{0}=p/q$, where $q$ is a positive integer and without loss of generality we will set $p=1$ throughout this paper. The enlarged unit cell is then obtained by extending the original unit cell $q$ times along the $y$ direction. This gives a magnetic BZ with $\pi\leq k_{x}a\leq\pi$, $\pi/q\leq k_{y}a\leq\pi/q$, $\pi\leq k_{z}a\leq\pi$. Working with the magnetic BZ will allow us to compute topological invariants and track the evolution of the Fermi arc states in a straightforward manner. Although $k_{x}$ in the magnetic BZ lies in the range $(-\pi,\pi)$ the spectrum along $k_{x}$ is in fact periodic with a period of $2\pi/q$ [54]. Therefore, in the analysis below we restrict ourselves to $k_{x}\in(-\pi/q,\pi/q)$.
Here we emphasize that, in contrast to Landau levels in the continuum, the Hofstadter Bloch bands arising at commensurate magnetic flux can host Weyl nodes and carry nonzero Chern numbers. Throughout this work, when we refer to a WSM phase in the presence of a magnetic field, we specifically mean that the corresponding Hofstadter Bloch bands exhibit Weyl nodes pinned at zero energy.
@@ IV.2 Phase diagram for $\gamma<0$
For the special case $\tilde{v}_{x}=v_{y}=1$, we can understand the transition between NI and LCI analytically by solving for the zeros of the Bloch-Hofstadter Hamiltonian in Eq. 9. These correspond to Weyl nodes in the magnetic BZ [42, 43], and their appearance is determined by the condition
@@
Figure 7: Energy gap as a function of separation between the Weyl nodes for sizable fluxes (a) $\phi_{B}/\phi_{0}=1/10$ and (b) $\phi_{B}/\phi_{0}=1/5$ for a fixed $\gamma=0.96$. See text for a detailed discussion of the sequence and structure of the transitions between the different gapped phases. (c) In a transition from an NI to an LCI the system goes through several closings and reopenings of the gap, as shown in Fig. 8. The red dots in (c) correspond to the values of $Q$ at which the energy bands in Fig. 8 are plotted.
@@
To understand the topological nature of the gapped phases, we calculate the $k_{x}$-dependent Chern number $C(k_{x})$ in the magnetic BZ. We find that for large $g$ (i.e. small $Q<Q^{*}$) $\sum_{n\in occ}C_{n}(k_{x})=0$ for all $k_{x}$, corresponding to an NI. On the other hand, for small $g$ (i.e. large $Q>Q^{*}$), $\sum_{n\in occ}C_{n}(k_{x})=1$ for all $k_{x}$, corresponding to an LCI phase. The transition between NI and LCI can be understood from the evolution of the solutions to Eq. (10). Let us fix the flux $1/q$ and look for solutions as $g$ (or equivalently $Q$) is varied. Starting from large $g$ (NI phase), when $g$ is decreased to $g_{2}$, the band gap closes and a Dirac point appears at the Gamma point of the magnetic BZ. Decreasing $g$ further leads to a splitting of the Dirac point into a pair of WNs that move away from each other towards the zone boundary. In this regime, solving in the slab geometry, a Fermi arc appears at the surface, and its length increases with the separation between the WNs. At a second critical value, $g=g_{1}$, the two WNs merge at the edge of the magnetic BZ $(\pm\pi/q,0,0)$ and disappear but the Fermi arc survives and spans the full magnetic BZ along $k_{x}$. Thus an LCI state emerges. This is demonstrated in Fig. 4.
@@ IV.3 Phase diagram for $\gamma>0$ and appearance of LCI
Figure 9: Evolution of zero energy surface states in the $k_{x}$-$k_{y}$ surface BZ in the transitions from an NI to an LCI state (the transition is depicted in Fig. 7(c)) for a fixed flux $1/5$ and $\gamma=0.96$. The $Q$ values are 2.175, 2.195, 2.22 and 2.30 in (a)-(d) respectively. Figure (a) shows the zero energy Fermi-arc states of the WSM state which appear in the transition NI to LCI. Chern numbers are calculated for a fixed $k_{x}$ value, indicated as vertical dashed lines representing ($k_{y},k_{z}$) planes over which we integrate the Berry curvature. (b) As $Q$ is increased, the Weyl nodes approach the edge of the magnetic BZ, merge, and get annihilated. At the end of this process, the corresponding Fermi arc spans the entire BZ along $k_{x}$. Thus an LCI state emerges. (c) A pair of Weyl nodes appear, accompanied by a new Fermi arc. This Fermi arc coexists with the surface states of the LCI bands. (d) As $Q$ is increased further, the Weyl nodes approach the edge of the magnetic BZ, merge at $(\pi/q,\pi/q)$, and get annihilated. Again, the corresponding Fermi arc now spans the entire BZ along $k_{x}$. Thus an LCI state emerges. In this state $C(k_{x})=0$ for all $k_{x}$.
@@ V Summary and Conclusion
An external orbital magnetic field, which is perpendicular to the direction of separation of Weyl cones of opposite chirality, introduces tunneling of electrons between the Weyl cones. This leads to a hybridization between the cones (note that such a hybridization is allowed since translation symmetry is broken by the presence of the orbital magnteic field) and an opening of a gap. Such a tunneling process requires overcoming the energy barrier separating the two Weyl nodes, leading to a nonperturbative dependence of the gap on the magnetic field. The induced gap is significant when the inverse magnetic length is comparable to or larger than the momentum-space separation between the Weyl nodes.
@@
We now turn to discuss briefly some of the experimental implications of our results. The magnetic field-induced gap opening in WSMs has direct consequences on the longitudinal magnetoresistance [41]. Such behavior has been already observed in experiments in the Weyl materials TaAs [36] and TaP [37]. However, longitudinal magnetoresistance does not give a direct indication of the topology of the gapped phases. In contrast, Hall conductivity measurements provide a clearer distinction between the gapped phases, namely the normal insulator, LCI, and LCI phases. A time-reversal symmetry-breaking Weyl semimetal with two Weyl nodes separated by a momentum space distance $Q$ along $k_{x}$ exhibits a non-quantized anomalous Hall conductivity of $\sigma_{yz}=\frac{e^{2}}{2\pi h}Q$ per layer along the $x$-direction. Upon entering the LCI phase with applied magnetic field, this Hall conductivity becomes quantized at $\sigma_{yz}=\frac{e^{2}}{h}$ per layer. In contrast, transitions to either a normal insulator or the LCI phase result in a vanishing Hall conductivity. To further distinguish between normal insulator and LCI, one may rely on surface-sensitive probes, such as STM, that can detect the presence of the zero-energy modes arising from the closed Fermi arc states in LCI.
We briefly comment on the effects of disorder, finite temperature, and interactions on the phase diagram. In the presence of disorder, the LCI phase cannot be sharply distinguished from a normal insulator, since the crystalline symmetry protecting the LCI phase is generically broken. Moreover, features associated with exponentially small gaps and with very narrow regions in parameter space are expected to be smeared out by disorder and thermal fluctuations. Nevertheless, the phase transitions from the WSM to either a normal insulating phase or an LCI phase are expected to remain robust, provided that disorder, temperature, and interaction scales are small compared to the magnetic-field-induced gap.
Finally, we comment on orbital magnetic fields that are not perfectly perpendicular to the separation of the Weyl nodes. The momentum parallel to the magnetic field, $k_{\parallel B}$, is a good quantum number. When the field has a component parallel to the separation between Weyl nodes, the value of $k_{\parallel B}$ differs at the two Weyl nodes, and they no longer hybridize directly. Hence, even a small misalignment of the field relative to the perpendicular direction can drastically suppress the gap opening [41]. However, this effect is mitigated in systems with nodal loops[55, 56, 57, 58, 59, 60, 61, 62]. In such systems, magnetic fields generically lead to direct hybridization of nodal states, regardless of the magnetic field orientation. This will be discussed in detail elsewhere.